40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 953     Entry time: Wed Sep 17 12:58:12 2008
Author: rob 
Type: Update 
Category: Locking 
Subject: bad 

Locking was pretty unsuccessful last night. All the subparts were locked (ARMs, PRM, DRM) and
aligned, but no DRMI+2ARMs locks. The alignment may have drifted significantly by the time I
got around to working the full shebang, however.

We should get back into the habit of clicking the
yellow "Restore last auto-alignment" button when we finish using the interferometer.
ELOG V3.1.3-